I have two more sisters than brothers. Each of my sisters has two more sisters than brothers. How many more sisters than brothers does my oldest brother have?
Answer by ikleyn(52786)

.

From the first two sentences, it is clear that "I", the person who tells the story, is one of sisters.


Let x be the number of boys in the family and y be the number of girls.

Them you have this equation


    y-1 = x + 2       (left side y-1 is the number of sisters to "I")


From this equation,


    y = x + 3.


It says that the number of girls in the family is 3 more than the number of boys.


Hence, the oldest brother (the oldest boy) has 4 more sisters than brothers.    ANSWER



For example the possible numbers of boys and girls can be


    boys      girls


     2         5

     3         6



Having this small table, you can check and confirm my final answer.

Solved.